# Cartesian subgroups in graph products of groups

### v1

@@ 6.4. Similarities with graph products of algebras
Our calculation of relations in the group $\mathrm{RC}_{\mathcal{K}}^{\prime}=\pi_{1}(\mathcal{R}_{\mathcal{K}})$ continues the list of parallel results between the case of “real” and “complex” moment-angle complexes $\mathcal{R}_{\mathcal{K}}=(D^{1},S^{0})^{\mathcal{K}}\simeq(\mathbb{R},\mathbb{R}\setminus\{0\})^{\mathcal{K}}$ and $\mathcal{Z}_{\mathcal{K}}:=(D^{2},S^{1})^{\mathcal{K}}\simeq(\mathbb{C},\mathbb{C}\setminus\{0\})^{\mathcal{K}}.$ Note that for flag complexes $\mathcal{R}_{\mathcal{K}}$ is aspherical (hence its homotopy properties are governed by the group $\pi_{1}(\mathcal{R}_{\mathcal{K}})$ theory), while $\mathcal{Z}_{\mathcal{K}}$ is simply connected and coformal [Vyl24, Corollary 6.8] (hence its homotopy properties are closely related to the associative algebra $H_{*}(\Omega\mathcal{Z}_{\mathcal{K}};\mathbf{k})$).
Many such results are discussed in [PR24, Introduction + Section 7]. Below we expand this list. (Here $\mathcal{K}$ is a flag simplicial complex on $[m],$ and $\mathbf{k}$ is a principal ideal domain.)
On the “homotopy theory side”, we have:
@@
    $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\text{ homogeneous generators modulo }\sum_{J\subset[m]}\mathrm{gen}\,H_{1}(\mathcal{K}_{J};\mathbf{k})\text{ homogeneous relations }$
    [Vyl24, Theorem 5.5] (or, as $\mathbb{Z}$-graded algebra, modulo $\sum_{n\geq 0}\mathrm{gen}\,\bigoplus_{|J|=n}H_{1}(\mathcal{K}_{J};\mathbf{k})$ homogeneous relations [Vyl24, Theorem 5.6]). This presentation is minimal.
@@
    $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\cdot\prod_{i\in J}\dim{\widetilde{H}}_{*}(\Omega X_{i};\mathbf{k})\text{ homogeneous generators modulo }\sum_{J\subset[m]}\widetilde{b}_{1}(\mathcal{K}_{J})\cdot\prod_{i\in J}\dim{\widetilde{H}}_{*}(\Omega X_{i};\mathbf{k})$
    homogeneous relations [Cai24, Corollary 5.14 + Proposition 4.4].
@@
    More generally, if $\underline{G}=(G_{1},\dots,G_{m})$ are finite groups, then the group $\operatorname{Cart}(\underline{G},\mathcal{K})=\pi_{1}((E\underline{G},\underline{G})^{\mathcal{K}})$ can be presented by
    $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\cdot\prod_{i\in J}\operatorname{rank}{\widetilde{H}}_{*}(\Omega BG_{i};\mathbb{Z})\text{ generators modulo }\sum_{J\subset[m]}\operatorname{rank}\Pi_{1}(\mathcal{K}_{J})\cdot\prod_{i\in J}\operatorname{rank}{\widetilde{H}}_{*}(\Omega BG_{i};\mathbb{Z})$
    relations by Theorem 1.1. (Here we use that $|G|-1=\operatorname{rank}{\widetilde{H}}_{*}(\Omega BG;\mathbb{Z})$.)

### v2

@@ 6.4. Similarities with graph products of algebras
Our calculation of relations in the group $\mathrm{RC}_{\mathcal{K}}^{\prime}=\pi_{1}(\mathcal{R}_{\mathcal{K}})$ continues the list of parallel results between the “real” and “complex” moment-angle complexes $\mathcal{R}_{\mathcal{K}}=(D^{1},S^{0})^{\mathcal{K}}\simeq(\mathbb{R},\mathbb{R}\setminus\{0\})^{\mathcal{K}}$ and $\mathcal{Z}_{\mathcal{K}}:=(D^{2},S^{1})^{\mathcal{K}}\simeq(\mathbb{C},\mathbb{C}\setminus\{0\})^{\mathcal{K}}.$ Many such results are discussed in [PR24, Introduction + Section 7]. Below we expand this list.
Here $\mathcal{K}$ is a flag simplicial complex on $[m],$ and $\mathbf{k}$ is a principal ideal domain. Note that for flag complexes $\mathcal{R}_{\mathcal{K}}$ is aspherical (hence its homotopy properties are governed by the group $\pi_{1}(\mathcal{R}_{\mathcal{K}})$), while $\mathcal{Z}_{\mathcal{K}}$ is simply connected and coformal [Vyl24, Corollary 6.8] (hence its homotopy properties are closely related to the associative algebra $H_{*}(\Omega\mathcal{Z}_{\mathcal{K}};\mathbf{k})$).
On the “associative algebra side”, we have:
@@
    $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\text{ generators modulo }\sum_{J\subset[m]}\mathrm{gen}\,H_{1}(\mathcal{K}_{J};\mathbf{k})\text{ relations}$
    [Vyl24, Theorem 5.5], where the generators and relations are homogeneous; this presentation is minimal among ×$\mathbb{Z}\times\mathbb{Z}_{\geq 0}^{m}$-homogeneous presentations. (If we consider $H_{*}(\Omega\mathcal{Z}_{\mathcal{K}};\mathbf{k})$ as a $\mathbb{Z}$-graded algebra, we actually need $\sum_{n\geq 0}\mathrm{gen}\,\bigoplus_{|J|=n}H_{1}(\mathcal{K}_{J};\mathbf{k})$ homogeneous relations [Vyl24, Theorem 5.6]).
@@
    $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\cdot\prod_{i\in J}\dim{\widetilde{H}}_{*}(\Omega X_{i};\mathbf{k})\text{ homogeneous generators modulo }$
    $\sum_{J\subset[m]}\widetilde{b}_{1}(\mathcal{K}_{J})\cdot\prod_{i\in J}\dim{\widetilde{H}}_{*}(\Omega X_{i};\mathbf{k})\text{ homogeneous relations}$
    [Cai24, Corollary 5.14 + Proposition 4.4].
@@
    More generally, if $\underline{G}=(G_{1},\dots,G_{m})$ are finite groups, then the group $\operatorname{Cart}(\underline{G},\mathcal{K})=\pi_{1}((E\underline{G},\underline{G})^{\mathcal{K}})=\pi_{1}((C\Omega B\underline{G},\Omega B\underline{G})^{\mathcal{K}})$ can be presented by
    $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\cdot\prod_{i\in J}\operatorname{rank}{\widetilde{H}}_{*}(\Omega BG_{i};\mathbb{Z})\text{ generators modulo }$
    $\sum_{J\subset[m]}\operatorname{rank}\Pi_{1}(\mathcal{K}_{J})\cdot\prod_{i\in J}\operatorname{rank}{\widetilde{H}}_{*}(\Omega BG_{i};\mathbb{Z})\text{ relations}$
    by Theorem 1.1. (Here we use that $|G|-1=\operatorname{rank}{\widetilde{H}}_{*}(\Omega BG;\mathbb{Z})$.)